OpenStudy (anonymous):

Find the interval on which f(x)= 1+x^(3/2)-(3/2)x - 1 is increasing so f ′′ (x)=3 /4x^(1/2)

OpenStudy (anonymous):

where do i go from there? is the answer -1, 0 or 0,1 or -1,1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Let me check your differentiation

OpenStudy (anonymous):

there is something wrong here

OpenStudy (anonymous):

x>1

OpenStudy (anonymous):

\[1+x^{\frac{3}{2}}-\frac{3}{2}x -1\] is this your equation?

OpenStudy (anonymous):

no one writes 1 - 1 so i am guessing there is a typo

OpenStudy (anonymous):

F''[x]= \[\frac{3}{4 \sqrt{x}}\]

OpenStudy (anonymous):

all u need to ensure is that you have a positive gradient. so differentiate once then set that greater than 0

OpenStudy (anonymous):

huh? lol :/

OpenStudy (anonymous):

can you please type out the actual function?

OpenStudy (anonymous):

you do not need the second derivative since you only need to know where it is increasing and decreasing.

OpenStudy (anonymous):

F''[x]=\[\frac{3x^{-\frac{1}{2}}}{4 }\] is never equal to to zero which means function is either increasing through out or decreasing throughout

OpenStudy (anonymous):

oh no! that is not right

OpenStudy (anonymous):

second derivative tells you concavity, not intervals of increasing and decreasing. i would still like to know what the function is

OpenStudy (anonymous):

All you need to do check the value of first derivative at any point to see whether it is positive or negative

OpenStudy (anonymous):

the original function is f(x)= 1+x^(3/2)-(3/2)x - 1

OpenStudy (anonymous):

the answer is x>1 man. u r confusing the poor girl

OpenStudy (anonymous):

does it really say 1 - 1?

OpenStudy (anonymous):

all those are possible answers that i got

OpenStudy (anonymous):

one of them is right the others are wrong

OpenStudy (anonymous):

\[f(x)=x^{\frac{3}{2}}-\frac{3}{2}x-1\]?

OpenStudy (anonymous):

you wrote \[f(x)=1+x^{\frac{3}{2}}-\frac{3}{2}x-1\] which makes little sense because no one writes plus one and minus one in the same function. i am thinking it is something else

OpenStudy (anonymous):

it could be (3/2)(x-1) satellite.but if nt then x>1 is the answer

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

f(x)= (1+x)^(3/2) - (3/2)x - 1

OpenStudy (anonymous):

sorry guys :/ silly error

OpenStudy (anonymous):

i don't know what the function is. @kboy you may be right, but how can you know since you don't know what the function is?

OpenStudy (anonymous):

ooooooooooh

OpenStudy (anonymous):

nw i get x>0

OpenStudy (anonymous):

find \[f'(x)=\frac{3}{2}\sqrt{x+1}-\frac{3}{2}\]

OpenStudy (anonymous):

so now you find the second derivative?

OpenStudy (anonymous):

set this equal 0 to find the critical points no

OpenStudy (anonymous):

Im so confused!!!

OpenStudy (anonymous):

you just need first derivative.

OpenStudy (anonymous):

\[\frac{3}{2}(\sqrt{x+1}-1)=0\] \[x=0\] is your only critical point

OpenStudy (anonymous):

you can do it that way if you like. i would not

OpenStudy (anonymous):

thus the interval is x>0 right??????????????

OpenStudy (anonymous):

if \[-1<x<0\] then this thing is negative and if \[x>1\] then it is positive. so original function was decreasing and then increasing.

OpenStudy (anonymous):

so increasing on \[(0,\infty)\]

OpenStudy (anonymous):

which is x>0..why u guys confuse the girl by making it sound so complicated

OpenStudy (anonymous):

lol. thanks to everyone though. I think I understand a little better now (:

OpenStudy (anonymous):

you just need the sign of the derivative. but you also need to know what the original function is which is what caused the confusion;.